Skip to end of metadata
Go to start of metadata

Another slow day, especially with rain that has come and is forecast to hang around for a while.  Most of the crew worked a half day.  Tensiometers still hadn't arrived by 6pm tonight, so we're anxious about tomorrow and the weekend.  I hope to update this shortly with news that they are here.  Otherwise, we'll have nothing to do until Monday...

However, due to DTU's efforts, we now have power to tse13.  Since this is a repeater, this allowed connection to several other towers that have been powered in the last few days.  Thus, we now have data from 6 towers streaming in.  At least one more tower is powered up and just needs a few WiFi settings to be adjusted.

In other news:

  • Finished prepping the valley sites (even the 3 towers that are not yet erected).  Now only have 1 more tower erect that hasn't been prepped.
  • Got the rudiments of wwwplots running.  This is being served up to the world at: http://ops-perdigao.dyndns.org/isf/projects/Perdigao/isfs/qcdata/  Scripts still need to be started to update this page automatically.
  • Apparently, Gordon has been updating cockpit, though we haven't started running it yet (due to no streams of data until late today).
  • The last (of 3) 4-component radiometer installations and its soil sensors are done.  We still have about 10 soils to install, but many are waiting for Samortecnica to finish work at these sites (since there has been a significant amount of surface disturbance as the towers are built).
  • Ted and Santiago were to have finished configuration of the project's WiFi gear.
  • Cornell leaves tomorrow, having installed their seismometers and 3 of 4 lidars.
  • DTU will have help from the municipality to install their 2 or 3 heavy lidars tomorrow.

 

  • No labels